TETSXPREME:

evaluate -(7^-3)(7^5)

TETSXPREME:

how do i figure this out?

HisKitten:

I'm not too sure, but I would try Distribution property.. thats what I would do.

Sailor:

7^5 is 16807 -7^-3 is -1/343 multiply those together and you'll have your answer

jhonyy9:

use this \[x^{-2} = \frac{ 1 }{ x^{2} }\] and (x^a) * (x^b) = x^(a+b)

Extrinix:

Remember that when multiplying with exponents, you add the exponents together, like Jhonyy is saying. \((x)^{a+b}\) Vice versa goes for division, in which you would subtract the exponents.

jhonyy9:

@extrinix wrote:
Remember that when multiplying with exponents, you add the exponents together, like Jhonyy is saying. \((x)^{a+b}\) Vice versa goes for division, in which you would subtract the exponents.
ATTENTION please !!! you will add the exponents just when the base powered numbers are the same !!!